x-intercept=4, m=2
Answer by ewatrrr(24785)

 
Hi
Using the standard slope-intercept form for an equation of a line y = mx + b
where m is the slope and b the y-intercept.
x-intercept=4, m=2
y = 2x + b |Pt(4,0) on the line: use ordered pair to solve for b
0 = 8 + b
-8 = b
y = 2x - 8